Find the volume generated by revolving the area bounded by the lines x+2y=4, x=0 and y=0, about x-axis.
1
Expert's answer
2020-12-21T18:36:48-0500
"x+2y=4=>y=-\dfrac{1}{2}x+2""y=0: 0=-\dfrac{1}{2}x+2=>x=4, Point (4,0)"
"V=\pi\displaystyle\int_{0}^4(-\dfrac{1}{2}x+2)^2dx=\pi\displaystyle\int_{0}^4(\dfrac{1}{4}x^2-2x+4)^2dx"

"=\pi\big[\dfrac{1}{12}x^3-x^2+4x\big]\begin{matrix}
 4 \\
 0
\end{matrix}=\pi(\dfrac{64}{12}-16+16-0)"

"=\dfrac{16\pi}{3} (units^3)"

"V=\dfrac{16\pi}{3}\ cubic\ units"